1. You need to make an aqueous solution of 0.138 M magnesium bromide for an experiment in lab, using a 125 mL volumetric flask. How much solid magnesium bromide should you add?

2. How many milliliters of an aqueous solution of 0.163 M aluminum nitrate is needed to obtain 13.4 grams of the salt?

3. n the laboratory you dissolve 19.8 g of calcium chloride in a volumetric flask and add water to a total volume of 250 . mL.
What is the molarity of the solution?
Answer #1

1)

volume , V = 1.25*10^2 mL

= 0.125 L

use:

number of mol,

n = Molarity * Volume

= 0.138*0.125

= 1.725*10^-2 mol

Molar mass of MgBr2,

MM = 1*MM(Mg) + 2*MM(Br)

= 1*24.31 + 2*79.9

= 184.11 g/mol

use:

mass of MgBr2,

m = number of mol * molar mass

= 1.725*10^-2 mol * 1.841*10^2 g/mol

= 3.176 g

Answer: 3.18 g

Only 1 question at a time please